# A determinantal formula for cluster variables in cluster algebras from surfaces

### v1

@@ 3 Snake graphs and cluster algebras
A tile is a graph isomorphic to a square drawn in the plane. A snake graph $\mathcal{G}$ is a plane graph obtained from a sequence of $d$ tiles $G_{1},\ldots,G_{d}$ such that each successive tile is placed on top or to the right of the previous tile, as in figure 1. That is, for $11\leq i\leq d-$ 1, tiles $G_{i}$ and $G_{i+1}$ share their right and left edge, respectively, or their top and bottom edge, respectively. The edges of $\mathcal{G}$ that lie on its unbounded face are called boundary edges, and the rest of the edges are internal. It is easily seen by induction that snake graphs are bipartite and always admit perfect matchings, and that a snake graph with $d$ tiles has $22d+$ 2 vertices and $13d+$ 1 edges.
Snake graphs were used in [8] as a crucial combinatorial tool to compute the cluster expansions of cluster variables in surface cluster algebras. Fix a bordered surface $(\Sigma,\mathbb{M})$ and an initial seed $(T,x_{0})$. As discussed in the introduction, each arc $\gamma$ in $\Sigma$ corresponds to a cluster variable $x_{\gamma}$, and if $\gamma\notin T$, then $x_{\gamma}$ is not part of the initial cluster $x_{0}$, so one must perform a series of mutations to find the expansion of $x_{\gamma}$ with respect to $x_{0}$. This can be avoided by virtue of the following.
@@
([8], Theorem 4.9) Orient $\gamma\notin T$ so that it crosses the edges $\tau_{i_{1}},\ldots,\tau_{i_{d}}$ of $T$ in order. Each $\tau_{i_{j}}$ is contained in exactly two triangles $\Delta_{j}$ and $\Delta_{j}^{\prime}$. Let $G_{j}$ be the quadrilateral formed by the triangles $\Delta_{j}$ and $\Delta_{j}^{\prime}$. Give each side of $G_{j}$ the weight $x_{k}$ if it is the internal diagonal of $T$ associated to $x_{k}$, and give it the weight 1 if it is a boundary edge of $T$. There is a weighted snake graph $\mathcal{G}_{\gamma}$ with tiles $G_{1},\ldots,G_{d}$ such that
$x_{\gamma}=\frac{1}{\operatorname{cross}(\gamma)}\sum_{P\in\operatorname{Match}(\mathcal{G}_{\gamma})}w(P)\,$
where $\operatorname{cross}(\gamma)=x_{i_{1}}x_{i_{2}}\cdots x_{i_{d}}$ and $w(P)$ is the product of the weights of the edges in $\mathcal{G}_{\gamma}$ that occur in $P$.
For more details on how the tiles $G_{j}$ are glued together to obtain $\mathcal{G}_{\gamma}$, see [8].
@@
By Theorem 2.2, snake graphs are pfaffian, and Theorem 2.1 suggests a way to produce a pfaffian orientation.
@@
Let $\mathcal{G}$ be snake graph with tiles $G_{1},G_{2},\ldots,G_{d}$. Define the orientation $\mathcal{P}(\mathcal{G})$ as follows:
@@
    Reverse the orientation of the left edge of $G_{1}$
The following is readily seen on account of Theorem 2.1. We give a direct proof for completeness.
@@
$\begin{pmatrix}z_{u_{1}w_{1}}&-z_{u_{2}w_{1}}\\ z_{u_{1}w_{2}}&z_{u_{2}w_{2}}\end{pmatrix}$
so that $\operatorname{pff}(\mathcal{P}(\mathcal{G}))=-z_{u_{1}w_{1}}z_{u_{2}w_{2}}-z_{u_{2}w_{1}}z_{u_{1}w_{2}}$. Take $\mathcal{G}$ with tiles $G_{1},\ldots,G_{d+1}$, $d\geq 1$ and give it the orientation $\mathcal{P}(\mathcal{G})$. Let $\mathcal{G}^{\prime}$ be the snake graph with tiles $G_{2},\ldots,G_{d+1}$ obtained by deleting the two leftmost vertices of $\mathcal{G}$. Observe that the induced orientation on $\mathcal{G}^{\prime}$ is $\mathcal{P}(\mathcal{G}^{\prime})$, so by induction the coefficients of the pfaffian of $\mathcal{G}^{\prime}$ all have the same sign $\epsilon\in\{1,-1\}$. By circling the boundary of $\mathcal{G}$ counterclockwise, we obtain a bipartition $\{u_{1},\ldots,u_{n}\}\cup\{w_{1},\ldots,w_{n}\}$ of its vertices. Label $V(\mathcal{G})$ so that $G_{1}$ has vertices $w_{n+1},u_{n},w_{n},u_{n+1}$ as in figure 1, and write $b_{i,j}$ and $b_{i,j}^{\prime}$ for the entries of $B_{\mathcal{P}(\mathcal{G})}(z)$ and $B_{\mathcal{P}(\mathcal{G}^{\prime})}(z)$, respectively. Note that $B_{\mathcal{P}(\mathcal{G})}(x)$ is obtained from $B_{\mathcal{P}(\mathcal{G}^{\prime})}(z)$ by adding two rows and two columns corresponding to vertices $u_{n+1}$ and $w_{n+1}$.
Any perfect matching $P$ of $\mathcal{G}$ either contains the edge $u_{n+1}w_{n+1}$ or the edges $u_{n+1}w_{n}$ and $u_{n}w_{n+1}$. In the first case, $P$ restricts to a perfect matching of $\mathcal{G}^{\prime}$, and its associated term in $\operatorname{pff}(\mathcal{P}(\mathcal{G}))$ is
$b_{P}$ $=\operatorname{sgn}\begin{pmatrix}1&2&\ldots&2n-1&2n&2n+1&2n+2\\ i_{1}&j_{1}&\ldots&i_{n}&j_{n}&2n+1&2n+2\end{pmatrix}b_{i_{1},j_{1}}\ldots b_{i_{n},j_{n}}(-z_{u_{n+1}w_{n+1}})$
$=\operatorname{sgn}\begin{pmatrix}1&2&\ldots&2n-1&2n\\ i_{1}&j_{1}&\ldots&i_{n}&j_{n}\end{pmatrix}b_{i_{1},j_{1}}\ldots b_{i_{n},j_{n}}(-z_{u_{n+1}w_{n+1}})$
@@
In the second case, $P$ contains edges $u_{n+1}w_{n}$ and $u_{n}w_{n+1}$. Such matchings are in bijective correspondence to matchings $P^{\prime}$ of $\mathcal{G}^{\prime}$ that contain the edge $u_{n}w_{n}$. The associated term in $\operatorname{pff}(\mathcal{P}(\mathcal{G}))$ is
$b_{P}$ $=\operatorname{sgn}\begin{pmatrix}1&2&\ldots&2n-1&2n&2n+1&2n+2\\ i_{1}&j_{1}&\ldots&2n-1&2n+2&2n+1&2n\end{pmatrix}b_{i_{1},j_{1}}\ldots b_{i_{n-1}j_{n-1}}(-z_{u_{n}w_{n+1}})(-z_{u_{n+1}w_{n}})\,.$
The corresponding term $b_{P^{\prime}}$ is
$\operatorname{sgn}\begin{pmatrix}1&2&\ldots&2n-1&2n\\ i_{1}&j_{1}&\ldots&2n-1&2n\end{pmatrix}b_{i_{1},j_{1}}\ldots b_{i_{n-1}j_{n-1}}z_{u_{n},w_{n}}.$
Note that $b_{P}$ can be obtained from $b_{P^{\prime}}$ by substituting $z_{u_{n},w_{n}}$ with $-z_{u_{n}w_{n+1}}z_{u_{n+1}w_{n}}$, since their associated permutations have opposite signs. Thus, $b_{P}=-\epsilon z_{u_{i_{1}},w_{j_{1}}}\cdots z_{u_{i_{n-1}},w_{j_{n-1}}}z_{u_{n}w_{n+1}}z_{u_{n+1}w_{n}}$, so every term of $\operatorname{pff}(\mathcal{P}(\mathcal{G}))$ has coefficient $-\epsilon$. ∎
Figure 2: An example in type $A$.
The previous proof also shows that the sign of each coefficient in $\operatorname{pff}(\mathcal{P}(\mathcal{G}))$ is $(-1)^{d}$, where $d$ is the number of tiles in $d$. Observe that the snake graph $T$ with one tile has pfaffian equal to $-\phi_{T}$. By induction, we see that $\operatorname{pff}(\mathcal{P}(\mathcal{G}))=(-1)^{d+1}\phi_{G}$.
@@
Let $\mathcal{G}$ be a snake graph with $d$ tiles, and let $M(z)$ be the weighted biadjacency matrix of $\mathcal{P}(\mathcal{G})$. Then
$\det M(z)=(-1)^{(d+1)(d+2)/2}\phi_{\mathcal{G}}\,.$
Proof.
$\mathcal{G}$ has $22d+$ 2 vertices, so $M(z)$ has size $d+1$. By Lemma 2.3,
$\det M(z)=(-1)^{d(d+1)/2}\operatorname{pff}(\mathcal{P}(\mathcal{G}))=(-1)^{d(d+1)/2+d+1}\phi_{\mathcal{G}}=(-1)^{(d+1)(d+2)/2}\phi_{\mathcal{G}}\,.$
Applying this corollary to Theorem 3.1, we obtain our determinantal formula.
Corollary 3.2.
Let $\mathcal{A}(\Sigma,\mathbb{M})$ be a surface cluster algebra with trivial coefficients, and let $(T,x_{0})$ be an initial seed. For every arc $\gamma\notin T$, let $\mathcal{G}_{\gamma}$ be its associated weighted snake graph, and $d$ be its number of tiles. Let $M$ be the matrix obtained by substituting each $z_{e}$ in $M_{\mathcal{P}(\mathcal{G}_{\gamma})}(z)$ with the weight of $e$ in $\mathcal{G}_{\gamma}$. The cluster expansion of $x_{\gamma}$ is given by
$x_{\gamma}=\frac{(-1)^{(d+1)(d+2)/2}}{\operatorname{cross}(\gamma)}\det M\,.$
@@
Consider the polygon in figure 2. The snake graph $\mathcal{G}$ associated to arc $\gamma$ (in red) is drawn to the right. In this case, $d=3$, so $(-1)^{(d+1)(d+2)/2}=1$. The matrix $M$ obtained from the weighted biadjacency matrix of $\mathcal{G}$ is
$M=\begin{pmatrix}1&0&0&x_{1}\\ -x_{2}&x_{4}&1&0\\ 0&-1&x_{6}&0\\ -x_{3}&0&-x_{5}&x_{4}\end{pmatrix}\,,$
and $\det(M)=x_{4}^{2}x_{6}+x_{4}+x_{1}x_{2}x_{5}+x_{1}x_{3}x_{4}x_{6}+x_{1}x_{3}$.
Acknowledgements
The author received support from Daniel Labardini Fragoso’s grant UNAM-PAPIIT-IN109824. This work was presented at the first meeting of the project ”cluster algebras and related topics” financed by bando Galileo 2024- G24-215 and held at the department S.B.A.I. of Sapienza-Università di Roma in June 2024. I thank the participants, in particular Giovanni Cerulli Irelli, Pierre-Guy Plamondon and Ralf Schiffler for valuable comments. The author is grateful to Giovanni Cerulli Irelli and Salvatore Stella for their hospitality.

### v2

@@ 3 Snake graphs and cluster algebras
A tile is a graph isomorphic to a square drawn in the plane. A snake graph $\mathcal{G}$ is a plane graph obtained from a sequence of $d$ tiles $G_{1},\ldots,G_{d}$ such that each successive tile is placed on top or to the right of the previous tile, as in figure 1. That is, for $1\leq i\leq d-1$, tiles $G_{i}$ and $G_{i+1}$ share their right and left edge, respectively, or their top and bottom edge, respectively. The edges of $\mathcal{G}$ that lie on its unbounded face are called boundary edges, and the rest of the edges are internal. It is easily seen by induction that snake graphs are bipartite and always admit perfect matchings, and that a snake graph with $d$ tiles has $2d+2$ vertices and $3d+1$ edges.
Snake graphs were used in [8] to provide manifestly positive expansions of the cluster variables in cluster algebras from surfaces. We briefly recall their construction. Fix a bordered surface $(\Sigma,\mathbb{M})$ and an initial seed $(\mathbf{x}_{0},\mathbf{y}_{0},B(T_{0}))$, where $T_{0}$ is an ideal triangulation and let $\mathcal{A}_{\bullet}(T_{0})$ be the associated cluster algebra with principal coefficients at 0. As discussed in the introduction, each tagged arc $\gamma$ in $\Sigma$ corresponds to a cluster variable $x_{\gamma}$, and if $\gamma\notin T_{0}$, then $x_{\gamma}$ is not part of the initial cluster $\textbf{x}_{0}$, so one must perform a series of mutations to find the expansion of $x_{\gamma}$ with respect to $\textbf{x}_{0}$. This can be avoided by virtue of the following.
@@
([8], Theorem 4.9) Orient $\gamma\notin T_{0}$ so that it crosses the diagonals $\tau_{i_{1}},\ldots,\tau_{i_{d}}$ of $T_{0}$ in order. Each $\tau_{i_{j}}$ is contained in exactly two triangles $\Delta_{j}$ and $\Delta_{j}^{\prime}$. Let $G_{j}$ be the quadrilateral formed by the triangles $\Delta_{j}$ and $\Delta_{j}^{\prime}$. Give each side $s$ of $G_{j}$ the weight $w(s)=x_{k}$ if it is the internal diagonal of $T_{0}$ associated to $x_{k}$, and give it the weight $w(s)=1$ if it is a boundary edge of $T_{0}$. There is a snake graph $\mathcal{G}_{\gamma}$ with tiles $G_{1},\ldots,G_{d}$ such that
$x_{\gamma}=\frac{1}{\operatorname{cross}(\gamma,T_{0})}\sum_{M\in\operatorname{Match}(\mathcal{G}_{\gamma})}x(M)y(M)\,$
where $x(M)$ is the product of all weights of the edges occurring in the matching $M$, $y(M)$ is the specialized height monomial of $M$, and $\operatorname{cross}(\gamma,T_{0}):=x_{i_{1}}x_{i_{2}}\cdots x_{i_{d}}$.
For complete details on how the tiles $G_{j}$ are glued together to obtain $\mathcal{G}_{\gamma}$, see [8]. There are exactly two matchings of $\mathcal{G}_{\gamma}$ consisting of boundary edges, $M_{-}$ and $M_{+}$. If tile $G_{1}$ is drawn in the plane in such a way that its orientation agrees with its orientation on $\Sigma$, $M_{-}$ is the matching that includes the south edge of $G$, and it is called the minimal matching of $\mathcal{G}_{\gamma}$. We recall the definition of the specialized height monomial from [9].
Definition 3.1.
Let $T_{0}=\{\tau_{1},\ldots,\tau_{n}\}$ be an ideal triangulation of $(\Sigma,\mathbb{M})$ and $\gamma\notin T_{0}$ a plain arc. By lemma 4.7 of [8], the symmetric difference $M\ominus M_{-}$ encloses a union of tiles $\cup_{j\in J}G_{i_{j}}$. The specialized height monomial of $M$ is given by
$y(M)=\prod_{j\in J}y_{\tau_{i_{j}}}\,.$
That is, $y(M)$ is the product of the variables $y_{i}$ that correspond to the diagonals of the tiles enclosed by $M\ominus M_{-}$. Note that several tiles of $\mathcal{G}_{\gamma}$ may have the same diagonal. By theorem 5.3 of [9], the set of perfect matchings of $\mathcal{G}_{\gamma}$ is a distributive lattice $\operatorname{Match(\mathcal{G}_{\gamma})}$. In this lattice, a matching $M$ covers another matching $M^{\prime}$ when $M$ can be obtained from $M^{\prime}$ via a local moved called a face twist: that is, there is a tile $G_{i}$ of $\mathcal{G}_{\gamma}$ with edges $e_{1},e_{2},e_{3},e_{4}$ such that $M\ominus M^{\prime}=\{e_{1},e_{2},e_{3},e_{4}\}$ and $M=(M^{\prime}\setminus\{e_{1},e_{2}\})\cup\{e_{3},e_{4}\}$. The perfect matching lattice of $\mathcal{G}_{\gamma}$ can be built up from sequences of face twists beginning at the minimal matching $M_{-}$. Thus, if the matching $M$ is obtained by applying the sequence of face twists at tiles $G_{j_{1}},\ldots,G_{j_{k}}$, $y(M)=y_{\tau_{j_{1}}}\cdots y_{\tau_{j_{k}}}$.
We now describe an alternative weight scheme for the edges of $\mathcal{G}_{\gamma}$ that recovers the specialized height monomial of each matching.
@@
(Principal weighting) Let $\mathcal{G}_{\gamma}$ be the snake graph on $n$ tiles associated to the arc $\gamma\notin T_{0}$ as in theorem 3.1, and let $w$ be its corresponding weighting. Consider the minimal matching $M_{-}$. We define the weighting $\tilde{w}$ of the edges of $\mathcal{G}_{\gamma}$ as follows: for each $1\leq i\leq n$
  • If it exists, pick a boundary edge of $e$ of $G_{i}$ such that $e\notin M_{-}$ and set $\tilde{w}(e)=y_{\tau_{i}}$.
  • If all boundary edges of $G_{i}$ are in $M_{-}$, then the consecutive tiles $G_{i-1},G_{i},G_{i+1}$ form a straight snake subgraph of $\mathcal{G}_{\gamma}$. Let $e,e^{\prime}$ be boundary edges of $G_{i-1}$ and $G_{i+1}$, respectively, and let $d,d^{\prime}$ be the edges of $G_{i}$ that are not in $M_{-}$. Set $\tilde{w}(e^{\prime})=\tilde{w}(d)=\tilde{w}(d^{\prime})=\tilde{w}(e^{\prime})=\sqrt{y_{i}}$.
  • For all other edges $a$ of $\mathcal{G}_{\gamma}$, set $\tilde{w}(a)=1$.
Let $w_{\bullet}$ be the weighting of $\mathcal{G}_{\gamma}$ defined by $w_{\bullet}(e):=w(e)\tilde{w}(e)$ for each edge $e$. We call $w_{\bullet}$ a principal weighting of $\mathcal{G}_{\gamma}$.
Proposition 3.1.
Let $\mathcal{G}_{\gamma}$ be the snake graph of the arc $\gamma\notin T_{0}$. For each $M\in\operatorname{Match}(\mathcal{G}_{\gamma})$, let $w_{\bullet}(M):=\prod_{e\in M}w_{\bullet}(e)$. Then
$x_{\gamma}=\sum_{M\in\operatorname{Match}(\mathcal{G}_{\gamma})}w_{\bullet}(M)\,.$
Proof.
By virtue of theorem 3.1, we just need to show that $w_{\bullet}(M)=x(M)y(M)$. By definition, $x(M)$ divides $w_{\bullet}(M)$ and there is no power of $x_{i}$ that divides $w_{\bullet}(M)/x(M)$, for any $i$. Thus, it is enough to show that each variable $y_{j}$ appears with equal multiplicity in both $w_{\bullet}(M)$ and $y(M)$. Now, $y_{j}$ appears once in $y(M)$ for each tile $G_{i}$ with diagonal $\tau_{j}$ that is enclosed by the edges $M\ominus M_{-}$. If there is a boundary edge $e$ of $G_{i}$ in $M$, then $e\notin M_{-}$, so $M$ includes all boundary edges of $G_{i}$ (one if $G_{i}$ is a corner tile, two otherwise), and therefore $y_{j}\mid w_{\bullet}(M)$. If there is no boundary edge of $G_{i}$ in $M$, then both are in $M_{-}$, and $G_{i}$ is in such a position as in the third tile of figure 2. Let $e,e^{\prime}$ be the edges in $G_{i}\cap M_{-}$, with $e$ adjacent to $a_{1},a_{2}$ and $e^{\prime}$ adjacent to $a_{1}^{\prime},a_{2}^{\prime}$. Since $G_{i}$ is enclosed by $\ominus M_{-}, MM$ must include one element $e_{1}$ of $\{a_{1},a_{2}\}$ and one element $e_{2}$ of $\{a_{1}^{\prime},a_{2}^{\prime}\}$. In all cases, $w_{\bullet}(e_{1})=\sqrt{y_{j}}=w_{\bullet}(e_{2})$, so $y_{j}\mid w_{\bullet}(M)$. Reciprocally, any matching $M$ that includes the edge $a_{1}$ must also include either $a_{1}^{\prime}$ or $a_{2}^{\prime}$, so $M\ominus M_{-}$ encloses $G_{i}$ and $y_{j}$ appears in both $y(M)$ and $w_{\bullet}(M)$. ∎
Now, by Theorem 2.2, snake graphs are pfaffian, and Theorem 2.1 immediately suggests a pfaffian orientation:
Definition 3.3.
Let $\mathcal{G}$ be a snake graph with tiles $G_{1},\ldots,G_{d}$. Define the orientation $\mathcal{P}(\mathcal{G})$ as follows:
@@
    Reverse the orientation of the left edge of $G_{1}$.
The following is readily seen on account of Theorem 2.1, but we give a direct proof.
@@
$\begin{pmatrix}z_{u_{1}w_{1}}&z_{u_{2}w_{1}}\\ -z_{u_{1}w_{2}}&z_{u_{2}w_{2}}\end{pmatrix}$
so that $\operatorname{pff}(\mathcal{P}(\mathcal{G}))=z_{u_{1}w_{1}}z_{u_{2}w_{2}}+z_{u_{2}w_{1}}z_{u_{1}w_{2}}$. Take $\mathcal{G}$ with tiles $G_{1},\ldots,G_{d+1}$, $d\geq 1$ and give it the orientation $\mathcal{P}(\mathcal{G})$. Let $\mathcal{G}^{\prime}$ be the snake graph with tiles $G_{2},\ldots,G_{d+1}$ obtained by deleting the two leftmost vertices of $\mathcal{G}$. Observe that the induced orientation on $\mathcal{G}^{\prime}$ is $\mathcal{P}(\mathcal{G}^{\prime})$, so by induction the coefficients of the pfaffian of $\mathcal{G}^{\prime}$ all have the same sign $\epsilon\in\{1,-1\}$. By circling the boundary of $\mathcal{G}$ counterclockwise, we obtain a bipartition $\{u_{1},\ldots,u_{n}\}\cup\{w_{1},\ldots,w_{n}\}$ of its vertices. Relabel $V(\mathcal{G})$ so that $G_{1}$ has vertices $w_{n+1},u_{n},w_{n},u_{n+1}$ as in figure 1, and write $a_{i,j}$ and $a_{i,j}^{\prime}$ for the entries of $A_{\mathcal{P}(\mathcal{G})}(z)$ and $A_{\mathcal{P}(\mathcal{G}^{\prime})}(z)$, respectively. Note that $A_{\mathcal{P}(\mathcal{G})}(x)$ is obtained from $A_{\mathcal{P}(\mathcal{G}^{\prime})}(z)$ by adding a row and a column corresponding to vertices $u_{n+1}$ and $w_{n+1}$.
Any perfect matching $M$ of $\mathcal{G}$ either contains the edge $u_{n+1}w_{n+1}$ or the edges $u_{n+1}w_{n}$ and $u_{n}w_{n+1}$. In the first case, $M$ restricts to a perfect matching of $\mathcal{G}^{\prime}$, and its associated term in $\operatorname{pff}(\mathcal{P}(\mathcal{G}))$ is
$a_{M}$ $=\operatorname{sgn}\begin{pmatrix}1&2&\ldots&2n-1&2n&2n+1&2n+2\\ i_{1}&j_{1}&\ldots&i_{n}&j_{n}&2n+1&2n+2\end{pmatrix}a_{i_{1},j_{1}}\ldots a_{i_{n},j_{n}}(-z_{u_{n+1}w_{n+1}})$
$=\operatorname{sgn}\begin{pmatrix}1&2&\ldots&2n-1&2n\\ i_{1}&j_{1}&\ldots&i_{n}&j_{n}\end{pmatrix}a_{i_{1},j_{1}}\ldots a_{i_{n},j_{n}}(-z_{u_{n+1}w_{n+1}})$
@@
In the second case, $M$ contains edges $u_{n+1}w_{n}$ and $u_{n}w_{n+1}$. Such matchings are in bijective correspondence to matchings $M^{\prime}$ of $\mathcal{G}^{\prime}$ containing $u_{n}w_{n}$. The associated term in $\operatorname{pff}(\mathcal{P}(\mathcal{G}))$ is
$a_{M}$ $=\operatorname{sgn}\begin{pmatrix}1&2&\ldots&2n-1&2n&2n+1&2n+2\\ i_{1}&j_{1}&\ldots&2n-1&2n+2&2n+1&2n\end{pmatrix}a_{i_{1},j_{1}}\ldots a_{i_{n-1}j_{n-1}}(-z_{u_{n}w_{n+1}})(-z_{u_{n+1}w_{n}})\,.$
The corresponding term $a_{M^{\prime}}$ is
$\operatorname{sgn}\begin{pmatrix}1&2&\ldots&2n-1&2n\\ i_{1}&j_{1}&\ldots&2n-1&2n\end{pmatrix}a_{i_{1},j_{1}}\ldots a_{i_{n-1}j_{n-1}}z_{u_{n},w_{n}}.$
Note $a_{M}$ can be obtained from $a_{M^{\prime}}$ by substituting $z_{u_{n},w_{n}}$ with $z_{u_{n}w_{n+1}}z_{u_{n+1}w_{n}}$, for their associated permutations have opposite signs. Thus, $a_{M}=-\epsilon z_{u_{i_{1}},w_{j_{1}}}\cdots z_{u_{i_{n-1}},w_{j_{n-1}}}z_{u_{n}w_{n+1}}z_{u_{n+1}w_{n}}$, so every term of $\operatorname{pff}(\mathcal{P}(\mathcal{G}))$ has coefficient $\epsilon$. ∎
Figure 2: Example in type AA. To the right, a principal weighting of $\mathcal{G}_{\gamma}$ with $M_{-}$ depicted in blue.
Remark 3.1.
As in the previous proof, one can label the vertices cyclically by circling the boundary of $\mathcal{G}$ in the clockwise direction, starting from the leftmost bottom vertex $v_{1}$. Call this the cyclical labeling of $\mathcal{G}$. With this labeling, the perfect matching of $\mathcal{P}(\mathcal{G})$ given by the boundary edges $v_{1}v_{2},v_{3}v_{4},\ldots,v_{2n-1}v_{2n}$ has positive sign in $\operatorname{pff}(\mathcal{G})$, as it is associated to the identity permutation. Thus, every coefficient of $\operatorname{pff}(\mathcal{P}(\mathcal{G}))$ is positive with respect to this labeling.
Remark 3.2.
The adjacency matrix of $\mathcal{G}$ with respect to the cyclic labeling is not of the form 2.1. To fix this, we relabel the vertices of $\mathcal{G}$ via the permutation $\sigma$ given by $\sigma(i)=(i+1)/2$ for odd $i$ and $\sigma(j)=i/2+n$ for even $j$, which maps the odd numbered vertices in the cyclic labeling of $\mathcal{G}$ onto the set $\{1,\ldots,n\}$, and the even numbered vertices onto the set $\{n+1,\ldots,2n\}$. Moreover, $\operatorname{sgn}(\sigma)=(-1)^{n(n-1)/2}$, as can be seen by counting the negative slope segments with endpoints in $\{(i,\sigma(i))\in\mathbb{R}^{2}\mid i\in\{1,\ldots,2n\}\}$.
Definition 3.4.
Let $\mathcal{G}$ be a snake graph. Let $B_{\mathcal{P}(\mathcal{G})}(z)$ denote the weighted biadjacency matrix of $\mathcal{G}$ obtained from the relabeling in remark 3.2. That is, when walking the boundary cycle of $\mathcal{G}$ in the counterclockwise sense starting from the southwest vertex of the first tile of $\mathcal{G}$, we get the sequence of labels $1,n/2+1,2,n/2+2,\ldots,n/2,n$.
Note that all non zero entries of $B_{\mathcal{P}(\mathcal{G})}(z)$ below the have negative sign below the diagonal, and positive sign above it. In view of lemma 2.3 and remarks 3.1 and 3.2, we have the following:
@@
Let $\mathcal{G}$ be a snake graph. Then $\det B_{\mathcal{P}(\mathcal{G})}(z)=\phi_{\mathcal{G}}$
Remark 3.3.
There is a shorter proof of proposition 3.2 that uses the lattice structure of $\operatorname{Match}(\mathcal{G})$. To wit, let the matching $M$ be covered by $M^{\prime}$, and write $\sigma_{M}$, $\sigma_{M^{\prime}}$ for their associated permutations, with respect to some labeling of $V(\mathcal{G})$. If $M^{\prime}$ is obtained by twisting the face $f$ with the vertices $u_{i},w_{j},u_{k},w_{l}$, then $\sigma^{\prime}=(i\ l)\sigma$, but since $f$ is oddly oriented by theorem 2.1, $b_{i,j}b_{l,k}=-b_{i,l}b_{j,k}$, so the terms corresponding to $M$ and $M^{\prime}$ have the same sign in $\operatorname{pff}\mathcal{G}$. Thus, we may more generally take any pfaffian orientation of $\mathcal{G}$ and label its vertices in such a way that for some given matching $M$ (say, the minimal matching $M_{-}$), the permutation $\sigma_{M}$ is the identity.
With remark 3.3 in mind, we give the following definition.
Definition 3.5.
Let $\mathcal{G}_{\gamma}$ be the snake graph associated to the plain arc $\gamma\notin T_{0}$. Fix any pfaffian orientation of $\mathcal{G}_{\gamma}$ and label its vertex set so that $\sigma_{M_{-}}=\operatorname{Id}$. Let $B(z)$ be the corresponding weighted biadjacency matrix of $\mathcal{G}_{\gamma}$. Denote by $B(w_{\bullet})$ the matrix obtained by substituting the weight $w_{\bullet}(e)$ in each entry of $B(z)$. We will say $B(w_{\bullet})$ is a principal biadjacency matrix of $\mathcal{G}_{\gamma}$.
We can now compactly state our determinantal formula, which follows from corollary 3.1.
Proposition 3.3.
Let $\mathcal{A}(\Sigma,\mathbb{M})$ be a surface cluster algebra with principal coefficients at the initial seed $(T_{0},\mathbf{x}_{0},\mathbf{y}_{0})$, where $T_{0}$ is an ideal triangulation of $\Sigma$. For every plain arc $\gamma\notin T_{0}$, let $\mathcal{G}_{\gamma}$ be its associated snake graph. Let $B(w_{\bullet})$ be a principal biadjacency matrix of $\mathcal{G}_{\gamma}$. Then
$x_{\gamma}=\frac{\det B(w_{\bullet})}{\operatorname{cross}(\gamma,T_{0})}\,.$
@@
Consider figure 2. Take the pfaffian orientation $\mathcal{P}(\mathcal{G}_{\gamma})$ and the depicted principal weighting. The corresponding principal biadjacency matrix is
$B=\begin{pmatrix}1&0&0&0&y_{2}\\ -x_{2}y_{3}&x_{3}&0&\sqrt{y_{4}}&0\\ 0&-x_{4}y_{6}&x_{5}&0&0\\ 0&-\sqrt{y_{4}}&-\sqrt{y_{4}}&x_{6}&0\\ -1&0&0&-x_{4}\sqrt{y_{4}}&x_{3}\end{pmatrix}\,$,
and $\det(B)=x_{3}x_{5}x_{6}y_{2}+x_{2}x_{4}^{2}y_{2}y_{3}y_{4}y_{6}+x_{2}x_{4}x_{5}y_{2}y_{3}y_{4}+x_{4}y_{2}y_{4}y_{6}+x_{5}y_{2}y_{4}+x_{3}^{2}x_{5}x_{6}+x_{3}x_{4}y_{4}y_{6}+x_{3}x_{5}y_{4}\,.$
Acknowledegments
The author received support from Daniel Labardini Fragoso’s grant UNAM-PAPIIT-IN109824. This work was presented at the first meeting of the project “Cluster algebras and related topics” financed by bando Galileo 2024- G24-215 and held at the department S.B.A.I. of Sapienza-Università di Roma in June 2024. The author thanks the participants, in particular Giovanni Cerulli Irelli, Pierre-Guy Plamondon and Ralf Schiffler for valuable comments, as well as Daniel Labardini-Fragoso for his kind support and encouragement. The author is grateful to Giovanni Cerulli Irelli and Salvatore Stella for their hospitality, and to Jon Wilson for his kind comments on a previous draft.